# Deep Bayes Factors

### v2

@@ 7 Real Data Analysis - Cognitive Bias
In psychological cognitive research, the use of Bayes factors to test among explanatory narratives has been on the rise (Lee and Wagenmakers, 2014; Heck et al., 2023). We study weapon identification task (Rivers, 2017), which delves into the impact of stereotypes on participants’ ability to recognize objects. Participants were shown images and asked to distinguish between tools and weapons. Each image presented either a tool or a weapon alongside one of three priming images: a white face, a black face, or a neutral face outline. The study assessed the accuracy of responses across all six combinations of primes (white, black, neutral) and targets (tool, weapon), with 36 trials for each combination, amounting to 216 trials per participant. We focus on the data subset where participants were given a 1,000 ms response deadline. This subset corresponds to a group of $N=42$ participants.
@@
Figure 9: Figure captured from Heck et al. (2023). Each branch of the probability tree either leads to a correct (+) or incorrect (-) response. The parameters refer to the probabilities that one of the hypothesized processes succeeds (C = controlled; A = automatic; B = guessing Tool).
@@
For this task, Bishara and Payne (2009); Heck et al. (2023) compared two multinomial-processing tree models: $M_{1}$, the process-dissociation model with guessing (PD), and $M_{2}$, the Stroop model with guessing (Stroop). Both models are depicted in Figure 9, each with a different order of cognitive processes. In essence, testing between the two models amounts to determining whether stereotypes (prejudice) are activated before a random guess in the recognition of weapons. In the diagrams, $A$ denotes the probability of automatic stereotype activation, and $B$ stands for the likelihood of a random guess. $C$ represents the probability that the controlled process is active. In the PD model, controlled processing probability $C$ is considered first. If controlled processing is not active (with a probability of $1-C$), an individual might rely on their automatic stereotype with probability $A$. If not, they might make a random guess with probability $B$. Conversely, in the Stroop model, the automatic stereotype is activated before the controlled process. In Bishara and Payne (2009); Heck et al. (2023), the total number of successes for each participant follows a binomial distribution, denoted as $Bin(36,p_{j})$, where $j=1,...,6$ represents the six prime-tool combinations. The corresponding values for $p_{j}$ under each model are presented in Table S3 in Section S5.
As a baseline we use TreeBUGS (Heck et al., 2018, 2023), an R package specifically crafted for hierarchical multinomial-processing tree modeling. Before the recent implementation of Bayes factor estimates inside this software, estimating Bayes factors for this type of modeling was challenging. Since TreeBUGS does not distinguish between individuals, in Heck et al. (2023), the data was processed into a single 12-dimensional vector (6 for success and 6 for failure count), i.e., condensed as $Bin(N\times 36,p_{j})$ for $j=1,...,6$. This approach inherently results in data always being represented as summary statistics.444 Otherwise, the model should be $Bin(36,p_{ij})$, with parameter $p_{ij}$ for each $i$-th participant. In our approach, while we apply our method to the 12-dimensional vector, we also consider the data as $N\times 6=252$ dimensions, leveraging DeepBF’s flexibility to treat each participant as an independent entity from $Bin(36,p_{i})$. We denote the results as DeepBF-sum and DeepBF-full, respectively. Since Section 6 finds that a BNN is better in larger $n$ and a FNN is better in small $n$, we use a BNN for the 252 dimensional case and a FNN for the 12 dimensional case. We also compute ABC estimates (Section S2) for both 252 and 12 dimensional vectors with the hyperparameters in Table S2, denoting the result as ABC-full and ABC-sum, respectively. Lastly, we can also compute GBF and ABF values by using our DeepBF techniques, as described in Section 2.3. For this, considering $(n_{x},n_{z})=(1,41)$, we train an additional FNN of input size $6$ and combined with DeepBF-full.
@@ 8 Conclusion
Now we mention a few future directions for this work. In this study, we did not focus on optimizing learning hyperparameters, including the network architecture. Our goal was not to fine-tune these aspects to peak performance. Hence, we kept hyperparameters such as the learning rate or the network designs the same across all datasets and settings. Future work could benefit from advanced tuning and more sophisticated network designs. Indeed, we think further exploration of neural network architecture is desirable. The current feed-forward networks do not guarantee order invariance, while Bayes factors are order invariant for i.i.d. models. Using set-invariant networks, like those in Zaheer et al. (2017) may provide a reasonable solution. As $n$ increases, the distribution of the true Bayes factors becomes simpler, while the model complexity concurrently rises. Addressing this paradox may involve attention-based set invariant networks (Lee et al., 2019) where the size of feature maps does not increase with the input dimension.
@@
Below we reformulate the assumptions from Kaji and Ročková (2022) in the context of marginal likelihood ratio estimation.
@@
for a non-negative sequence $\delta_{T,n}$ where $\mathbb{M}_{T}(\cdot)$ was defined in (6). Moreover, assume that the bracketing entropy integral (defined in Section 1 in Appendix of Kaji and Ročková (2022)) satiesfies $J_{[]}(\delta_{T,n},\mathcal{D}_{n}^{*}(\delta_{T,n}),d_{H})\lesssim\delta_{T,n}^{2}\sqrt{T}$ and that there exists $\alpha<2$ such that $J_{[]}(\delta,\mathcal{D}_{n}^{*},d_{H})/\delta^{\alpha}$ has a majorant decreasing in $\delta$.
Under this assumption Kaji et al. (2020) conclude that $d_{H}(\hat{D}_{T},D^{*})=O_{\bar{P}}(\delta_{T,n})$, where $\bar{P}$ encompasses all randomness in the ABC reference table (simulated data from models $M_{1}$ and $M_{2}$). The next assumption is needed to show consistency of the Bayes factor estimator and can be regarded as a version of a bounded marginal likelihood condition. More details can be found in Kaji and Ročková (2022).
@@
Following the same steps as in the proof of Theorem 4.1 in Kaji and Ročková (2022) with notation $P_{0}=E_{1}$ and $D_{\theta}=D^{*}$ it can be verified [^5] Their proof presents a variant of the results with the expectation inside the absolute value. Inspecting the proof, however, it follows that these statements hold with the expectation outside the absolute value as well. that for any $D\in\mathcal{D}_{n}^{*}(\delta_{T,n})$
@@
For any $\delta>0$, denote by $D_{n}(\delta)=\{Y_{0}^{(n)}:n^{d_{1}/2}r(Y_{0}^{(n)})>1/\delta\}$ then it follows from the proof of Theorem 4.1 in Tokdar et al. (2010) that under Assumption 1
@@
and where with $\hat{D}_{T}$ defined in (5) and $D^{*}$ in (4). Next, we follow the steps of Theorem 4.1 in Tokdar et al. (2010). We utilize the following identity valid for any set $A_{n}^{1}\subset\mathbb{R}^{d_{1}}$ of parameters $\theta_{1}$ and $Y\in\mathcal{Y}^{n}$
@@
Next, under the model $M_{2}$ we can write $p^{*}(\cdot)=\pi_{2}(\cdot\,|\>\theta^{*}_{2})$ for some $\theta_{2}^{*}\in\mathbb{R}^{d_{2}}$. Similarly as in the proof of Theorem 4.1 in Tokdar et al. (2010), Assumption 1 then implies that for any $\delta>0$
@@
Given a data sample $Y_{0}$, the widely used ABC algorithm for Bayes factor estimation for $BF_{1,2}(Y_{0})$ accepts an ABC sample $\tilde{Y}$ if $\rho(\eta(Y_{0}),\eta(\tilde{Y}))\leq\epsilon$, where $\eta$ is a sufficient statistic, $\rho$ is a distance measure, and $\epsilon$ is the threshold. Choosing the right $\epsilon$ is a particularly challenging task. A too large $\epsilon$ reduces the quality of the estimator, while a too small $\epsilon$ increases computational burdens. To accept a pre-fixed number of samples, many more simulated samples will be required. Therefore, to step side the difficulty of tuning $\epsilon$, we use a rank-based acceptance rule. We first simulate $M$ samples $A=\{\tilde{Y}_{1},...,\tilde{Y}_{M}\}$, and then compute the distances for all $Y_{i}$ as $\rho(\eta(Y_{0}),\eta(\tilde{Y}_{i}))$. Lastly, we take the samples of the $k$-smallest distances. Other than this new acceptance rule, this rank-based ABC estimator is identical to the ABC-based estimator explained in Robert et al. (2011).
@@
Throughout the paper, we employed a fully connected feed-forward neural network (FNN), renowned for its universal approximability (Hornik, 1991). An FNN consists of a linear layer followed by a non-linear activation, which in our case is ReLU. We observed that expanding intermediate layers proportionally with input size can improve estimation but also introduce optimization instability. To address this, we utilized batch normalization neural networks (BNNs) (Ioffe and Szegedy, 2015). In BNNs, each linear layer is followed by batch normalization before applying ReLU, except for the output layer, which uses a sigmoid activation. The complete design of our neural network is detailed in Section S3.1. For FNNs, the architecture includes an input layer mapping to a hidden space of size 64, followed by two $64\times 64$ linear layers with ReLU activations, and a final linear layer that maps to a single output with a sigmoid activation. Conversely, for BNNs, we upscale the first hidden layer based on the input size to improve estimation. Note that BNNs might show inconsistencies between training and testing due to different mean and variance handling. Therefore, during evaluation, we use normalization computed from a single mini-batch. We trained our models using the Adam optimizer (Kingma and Ba, 2014) with default hyperparameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$), an initial learning rate $0.01$, and a learning rate decay of 0.99 every 1000 iterations. The training loss function did not include regularization, and the networks were trained without dropout for $M=400,000$ iterations using a mini-batch size of 200, resulting in a total of $T=80,000,000$ samples.

### v3

@@ 7 Real Data Analysis - Cognitive Bias
In psychological cognitive research, the use of Bayes factors to test among explanatory narratives has been on the rise [47, 33]. We study weapon identification task [65], which delves into the impact of stereotypes on participants’ ability to recognize objects. Participants were shown images and asked to distinguish between tools and weapons. Each image presented either a tool or a weapon alongside one of three priming images: a white face, a black face, or a neutral face outline. The study assessed the accuracy of responses across all six combinations of primes (white, black, neutral) and targets (tool, weapon), with 36 trials for each combination, amounting to 216 trials per participant. We focus on the data subset where participants were given a 1,000 ms response deadline. This subset corresponds to a group of $N=42$ participants.
@@
Figure 7: Figure captured from [33]. Each branch of the probability tree either leads to a correct (+) or incorrect (-) response. The parameters refer to the probabilities that one of the hypothesized processes succeeds (C = controlled; A = automatic; B = guessing Tool).
@@
For this task, [12, 33] compared two multinomial-processing tree models: $M_{1}$, the process-dissociation model with guessing (PD), and $M_{2}$, the Stroop model with guessing (Stroop). Both models are depicted in Figure 7, each with a different order of cognitive processes. In essence, testing between the two models amounts to determining whether stereotypes (prejudice) are activated before a random guess in the recognition of weapons. In the diagrams, $A$ denotes the probability of automatic stereotype activation, and $B$ stands for the likelihood of a random guess. $C$ represents the probability that the controlled process is active. In the PD model, controlled processing probability $C$ is considered first. If controlled processing is not active (with a probability of $1-C$), an individual might rely on their automatic stereotype with probability $A$. If not, they might make a random guess with probability $B$. Conversely, in the Stroop model, the automatic stereotype is activated before the controlled process. In [12, 33], the total number of successes for each participant follows a binomial distribution, denoted as $Bin(36,p_{j})$, where $j=1,...,6$ represents the six prime-tool combinations. The corresponding values for $p_{j}$ under each model are presented in Table S3 in Section S7.
@@
As a baseline we use TreeBUGS [32, 33], an R package specifically crafted for hierarchical multinomial-processing tree modeling. Before the recent implementation of Bayes factor estimates inside this software, estimating Bayes factors for this type of modeling was challenging. Since TreeBUGS does not distinguish between individuals, in [33], the data was processed into a single 12-dimensional vector (6 for success and 6 for failure count), i.e., condensed as $Bin(N\times 36,p_{j})$ for $j=1,...,6$. This approach inherently results in data always being represented as summary statistics.444 Otherwise, the model should be $Bin(36,p_{ij})$, with parameter $p_{ij}$ for each $i$-th participant. In our approach, while we apply our method to the 12-dimensional vector, we also consider the data as $N\times 6=252$ dimensions, leveraging DeepBF’s flexibility to treat each participant as an independent entity from $Bin(36,p_{i})$. We denote the results as DeepBF-sum and DeepBF-full, respectively. Since Section 6 finds that a BNN is better in larger $n$ and a FNN is better in small $n$, we use a BNN for the 252 dimensional case and a FNN for the 12 dimensional case. We also compute ABC estimates (Section S2) for both 252 and 12 dimensional vectors with the hyperparameters in Table S2, denoting the result as ABC-full and ABC-sum, respectively. Lastly, we can also compute GBF and ABF values by using our DeepBF techniques, as described in Section 2.3. For this, considering $(n_{x},n_{z})=(1,41)$, we train an additional FNN of input size $6$ and combined with DeepBF-full.
@@ 8 Conclusion
Now we mention a few future directions for this work. In this study, we did not focus on optimizing learning hyperparameters, including the network architecture. Our goal was not to fine-tune these aspects to peak performance. Hence, we kept hyperparameters such as the learning rate or the network designs the same across all datasets and settings. Future work could benefit from advanced tuning and more sophisticated network designs. Indeed, we think further exploration of neural network architecture is desirable. The current feed-forward networks do not guarantee order invariance, while Bayes factors are order invariant for i.i.d. models. Using set-invariant networks, like DeepSets in [92] may provide a reasonable solution. As $n$ increases, the distribution of the true Bayes factors becomes simpler, while the model complexity concurrently rises. Addressing this paradox may involve attention-based set invariant networks [46] where the size of feature maps does not increase with the input dimension. A preliminary investigation of this idea using DeepSets is presented in Section S4 of the Supplement, which demonstrates promising results.
@@
Below we reformulate the assumptions from [41] in the context of marginal likelihood ratio estimation.
@@
for a non-negative sequence $\delta_{T,n}$ where $\mathbb{M}_{T}(\cdot)$ was defined in (6). Moreover, assume that the bracketing entropy integral (defined in Section 1 in Appendix of [41]) satiesfies $J_{[]}(\delta_{T,n},\mathcal{D}_{n}^{*}(\delta_{T,n}),d_{H})\lesssim\delta_{T,n}^{2}\sqrt{T}$ and that there exists $\alpha<2$ such that $J_{[]}(\delta,\mathcal{D}_{n}^{*},d_{H})/\delta^{\alpha}$ has a majorant decreasing in $\delta$.
Under this assumption [42] conclude that $d_{H}(\hat{D}_{T},D^{*})=O_{\bar{P}}(\delta_{T,n})$, where $\bar{P}$ encompasses all randomness in the ABC reference table (simulated data from models $M_{1}$ and $M_{2}$). The next assumption is needed to show consistency of the Bayes factor estimator and can be regarded as a version of a bounded marginal likelihood condition. More details can be found in [41].
@@
Following the same steps as in the proof of Theorem 4.1 in [41] with notation $P_{0}=E_{1}$ and $D_{\theta}=D^{*}$ it can be verified [^5] Their proof presents a variant of the results with the expectation inside the absolute value. Inspecting the proof, however, it follows that these statements hold with the expectation outside the absolute value as well. that for any $D\in\mathcal{D}_{n}^{*}(\delta_{T,n})$
@@
For any $\delta>0$, denote by $D_{n}(\delta)=\{Y_{0}^{(n)}:n^{d_{1}/2}r(Y_{0}^{(n)})>1/\delta\}$ then it follows from the proof of Theorem 4.1 in [83] that under Assumption 1 in the main text
@@
and where with $\hat{D}_{T}$ defined in (5) and $D^{*}$ in (4). Next, we follow the steps of Theorem 4.1 in [83]. We utilize the following identity valid for any set $A_{n}^{1}\subset\mathbb{R}^{d_{1}}$ of parameters $\theta_{1}$ and $Y\in\mathcal{Y}^{n}$
@@
Next, under the model $M_{2}$ we can write $p^{*}(\cdot)=\pi_{2}(\cdot\,|\>\theta^{*}_{2})$ for some $\theta_{2}^{*}\in\mathbb{R}^{d_{2}}$. Similarly as in the proof of Theorem 4.1 in [83], Assumption 1 then implies that for any $\delta>0$
@@
Given a data sample $Y_{0}$, the widely used ABC algorithm for Bayes factor estimation for $BF_{1,2}(Y_{0})$ accepts an ABC sample $\tilde{Y}$ if $\rho(\eta(Y_{0}),\eta(\tilde{Y}))\leq\epsilon$, where $\eta$ is a sufficient statistic, $\rho$ is a distance measure, and $\epsilon$ is the threshold. Choosing the right $\epsilon$ is a particularly challenging task. A too large $\epsilon$ reduces the quality of the estimator, while a too small $\epsilon$ increases computational burdens. To accept a pre-fixed number of samples, many more simulated samples will be required. Therefore, to step side the difficulty of tuning $\epsilon$, we use a rank-based acceptance rule. We first simulate $M$ samples $A=\{\tilde{Y}_{1},...,\tilde{Y}_{M}\}$, and then compute the distances for all $Y_{i}$ as $\rho(\eta(Y_{0}),\eta(\tilde{Y}_{i}))$. Lastly, we take the samples of the $k$-smallest distances. Other than this new acceptance rule, this rank-based ABC estimator is identical to the ABC-based estimator explained in [66].
@@
Throughout the paper, we employed a fully connected feed-forward neural network (FNN), renowned for its universal approximability [34]. An FNN consists of a linear layer followed by a non-linear activation, which in our case is ReLU. We observed that expanding intermediate layers proportionally with input size can improve estimation but also introduce optimization instability. To address this, we utilized batch normalization neural networks (BNNs) [36]. In BNNs, each linear layer is followed by batch normalization before applying ReLU, except for the output layer, which uses a sigmoid activation. The complete design of our neural network is detailed in Section S3.1. For FNNs, the architecture includes an input layer mapping to a hidden space of size 64, followed by two $64\times 64$ linear layers with ReLU activations, and a final linear layer that maps to a single output with a sigmoid activation. Conversely, for BNNs, we upscale the first hidden layer based on the input size to improve estimation. Note that BNNs might show inconsistencies between training and testing due to different mean and variance handling. Therefore, during evaluation, we use normalization computed from a single mini-batch. We trained our models using the Adam optimizer [45] with default hyperparameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$), an initial learning rate $0.01$, and a learning rate decay of 0.99 every 1000 iterations. The training loss function did not include regularization, and the networks were trained without dropout for $M=400,000$ iterations using a mini-batch size of 200, resulting in a total of $T=80,000,000$ samples.
@@
Appendix S4 DeepSet for Deep Bayes Factors
In this section, we apply the DeepSet network in [92] as a $q$-dimensional feature extractor before feeding it into an FNN. In this case, the input dimension for the FNN is $q$ (set as $q=2$ in our implementation). The DeepSet and FNN together compose the discriminator denoted by $D^{(t)}$ in Algorithm 1. The structure of the DeepSet is shown at the end of this section (Section S4.1). The benefit of this structure is that the DeepSet feature extractor makes the entire $D^{(t)}$ order-invariant among observations, and the size of the network does not scale up as $n$ increases.
For the setting in Example 1, we draw a figure corresponding to Figure 1 and Figure S18 with the automatic sufficient statistics learning using DeepSet, for increasing sufficient statistics dimensions in Figure S10 (high probability area) and Figure S11 (wider range). As we can see, within the high probability area, the estimation is reasonable, closely matching the quality of FNNs for $n=2$.
Figure S10: True and estimated BF’s on the $\log_{10}$ scale (DeepBF with DeepSet) evaluated on a fine grid of values $Y=(Y_{1},Y_{2})^{\prime}$. First row: Negative Binomial vs Poisson distributions. Second row: Gaussian vs Gaussian mixture priors. Third row: Hierarchical exponential distribution.
Figure S11: The scatter plot of random 30,000 samples on the visualization in Figure S10.
We also apply the DeepSet design to the settings considered in Section 6.1 (Dataset 1 (i) and Dataset 1 (ii)) and Section S6 (Dataset 3). We observed that the discriminator tends to collapse to a constant function as $n$ increases. This issue can be mitigated by using a larger DeepSet network architecture, which can be achieved by stacking multiple DeepSet layers, as described in Section S4.1. In this simulation, we stacked up two DeepSet layers. To further prevent collapse, we reduced the training mini-batch size to 300. Various alternative approaches to stabilize learning could also be considered, such as gradient clipping. To know the success of learning, we can use the classification loss in (6) or classification accuracy as a proxy. This allows us to select the best model from multiple runs with different random initializations.
The results of this approach are presented in Figure S12. The reported model was selected based on its highest classification accuracy on a test dataset (size $3000$) among 10 random runs. The benefits of using DeepSet for DeepBF extend beyond simply satisfying order invariance. This design enables the DeepBF estimator to handle much larger values of $n$. While we observe a decline in estimation quality as $n$ increases, the inferential consistency remains intact. Our theoretical results and proofs suggest that increasing the capacity of the DeepSet network—enabling it to represent an ideal classifier—could further enhance estimation quality. However, this would require longer training durations and the development of more stable learning techniques and hyperparameter configurations to reliably train larger networks.
Figure S12: The scatterplot of random 3,000 samples by DeepBF with the DeepSet architecture. It scales to much larger $n$ than the FNNs and BNNs (both computational time and learnability). In the plot, “cor” indicates the correlation coefficient between the true and estimated Bayes factor values.
S4.1 DeepSet Network architecture
The DeepSet architecture for DeepBF is “DeepSet2” which uses two “basic_DeepSet” layers. On the output of “DeepSet2”, a Net (FNN) is applied to get the final one dimensional output.
import torch
from torch import nn
class basic_DeepSet(nn.Module):
def __init__(self, input_dim, ss_dim,
c_factor = 64,c_n_layers=3,
device=”cuda”):
super(basic_DeepSet,self).__init__()
self.common_feature_net = MLP(device=device,
dim=input_dim,
z_dim = ss_dim,
dropout=0,
factor=c_factor,
n_layers=c_n_layers)
self.to(device)
self.device = device
def forward(self, x):
shape = x.shape
assert len(shape)==3
phi = self.common_feature_net(x.view(-1,shape[-1])).view(x.shape[0],x.shape[1],-1).mean(1)
return phi
class DeepSet2(Net):
def __init__(self, N_y, ss_dim=2, factor=64, depth=2, ss_dim_med=16, device=”cuda”):
super(DeepSet2, self).__init__(ss_dim, factor, depth)
self.suff_net1 = basic_DeepSet(device=device,
input_dim = 1,
ss_dim = ss_dim_med)
self.suff_net2 = basic_DeepSet(device=device,
input_dim = 1,
ss_dim = ss_dim)
def forward(self,x):
med = self.suff_net1(x).unsqueeze(-1)
med2= self.suff_net2(med)
return super().forward(med2)
class Net(nn.Module):
def __init__(self,input_shape,factor=64, depth = 2):
super(Net,self).__init__()
self.fc1 = nn.Linear(input_shape,factor)
self.linears = nn.ModuleList([nn.Linear(factor,factor) for i in range(depth)])
self.fc4 = nn.Linear(factor,1)
def forward(self,x):
x = torch.relu(self.fc1(x))
for md in self.linears:
x = torch.relu(md(x))
x = torch.sigmoid(self.fc4(x))
return x
Appendix S5 Examples
S5.1 Details of Example 1
All of these examples assume an iid data setup. Data 1 explores classical Bayesian modeling with Negative Binomial and Poisson distributions [11]. Define $M_{1}$ through $Y_{i}|p\sim NB(1,p)$, where $p\sim Beta(\alpha_{1},\beta_{1})$, and $M_{2}$ with $Y_{i}|\lambda\sim Pois(\lambda)$ where $\lambda\sim Gamma(\alpha_{2},\beta_{2})$. We set $\alpha_{1}=\beta_{1}=\alpha_{2}=\beta_{2}=1$. In this case, [66] identified bias in ABC Bayes factors (2) when a summary statistic is used. Data 2 considers Gaussian mixture data with a Gaussian prior on the means. Namely, $M_{1}:Y_{i}\sim N(\mu_{11},2^{2})/2+N(\mu_{12},2^{2})/2$ with $\mu_{11}\sim N(2,1.5^{2}),~{}\mu_{12}\sim\frac{1}{2}N(-2,1.5^{2})$, and $M_{2}:Y_{i}\sim N(\mu_{2},2.5^{2})$ with $\mu_{2}\sim N(0,1)$. Data 3 presents a nested model test, distinguishing between a hierarchical Bayesian model $M_{1}:Y_{i}|\lambda\sim Exp(\lambda)$ with $\lambda\sim Gamma(2,2)$ and a fixed-parameter model $M_{2}:Y_{i}\sim Exp(3)$. We can conceptualize $M_{2}$ as a Bayesian hierarchical model with a Dirac prior $\delta(\lambda-3)$. To nicely visualize heatmaps of the results, we start with two observations, i.e. $n=2$, and present results for larger $n$ later in Section 6.1. The heatmaps zoom in onto values $Y=(Y_{1},Y_{2})^{\prime}$ that are most likely under either one of the two considered models (See Figure S18 in Section S7 for a broader range of $Y$). The ABC implementation is detailed in Section S2. In the effort to see the best ABC performance, we intentionally use the full data $Y$ as a summary statistic. For the deep neural network architecture, we use an FNN. Details on the implementations and deep learning architectures are in Section S3.
S5.2 Toy Example for Model Criticism
Figure S13: Model Criticism on Data 1. The empirical density (KDE) plot of $Z(\tilde{Y},\hat{d})$ values defined in (12). Red lines indicate the $2.5\%$ and $97.5\%$ quantiles of the simulated $Z(\tilde{Y},\hat{d})$ values. The blue line represents $0.5$.
Example 3 (Toy Example).
Using data setups from Example 1 and 2, we train Bayesian GAN posterior generators [89]. For each dataset, we generate one fixed $Y_{0}^{(n)}\sim M_{j}$ and then simulate $\tilde{Y}$ according to $\pi(\cdot\,|\>Y_{0}^{(n)},M_{k})$ for four model combinations $(j,k)\in\{1,2\}\times\{1,2\}$ and $n=128$. For $M_{2}$ of Data 3, we instead sample $\tilde{Y}\sim\pi(\cdot\,|\>M_{2})$ independently from $Y_{0}$. [^6] The model parameter is fixed as opposed to random with a prior. Additionally, we manually set ”outliers” as values $Y_{0}^{(n)}$ unlikely under the given model777 All outliers are in a form of $Y_{\rm out}=(\alpha,\alpha,\cdots,\alpha)\in\mathcal{Y}^{n}$ for a manually chosen $\alpha$. Dataset 1: $\alpha=18$, Dataset 2: $\alpha=10$, Dataset 3: $\alpha=2.5$.. In Figure 1, the first two dimensions of the outliers (red) and typical inliers (yellow) are marked as stars. To simulate $Z(\tilde{Y},\hat{d})$ values defined in (12), we train a simple logistic regression classifier, where the non-linear predictor is a quadratic function of a univariate input. The classifiers are optimized for every new draw of $\tilde{Y}$ and evaluated on a new draw of $\tilde{Y}$ to compute the draw $Z(\tilde{Y},\hat{d})$. We had 3 00030003\,0003 000 draws for each model.
The kernel density smoothed empirical distribution of $Z(\tilde{Y},\hat{d})$ is displayed in Figure S13 for Data 1, and in Figure S22 and S23 (Section S7) for Data 2 and 3, respectively. The consonant cases $\tilde{Y}\sim\pi(\cdot\,|\>Y_{0}^{(n)},M_{k})$ and $Y^{(n)}_{0}\sim\pi(\cdot\,|\>M_{k})$ for $k=1,2$ are illustrated on the diagonal of the Figure S13 (panels (a) and (e)). We can see that when the posterior predictive distribution matches the true generating distribution of $Y_{0}^{(n)}$, the $Z(\tilde{Y},\hat{d})$ values are concentrated near $0.5$. On the other hand, panels (b) and (d) show the dissonant cases where the true generating model and predictive model are mismatched. In these plots, the $Z(\tilde{Y},\hat{d})$ values are more spread out and the value $0.5$ is barely included within the the $2.5\%$ and $97.5\%$ sample quantiles. This indicates poor or borderline evidence for the considered model suggesting further model comparisons through Bayes factors. On the other hand, when $Y_{0}^{(n)}$ are implausible under the considered model, critiquing the suggested model becomes easier. This is evident from panels (c) and (f) where the kernel density plots of sampled values $Z(\tilde{Y},\hat{d})$ are far away from $0.5$. Such a scenario yields a more decisive rejection of the proposed model without the need for further Bayes factor model comparisons. It is important to note, however, Bayes factor values evaluated at observed data $Y^{(n)}_{0}$ implausible under both models may still strongly prefer one model over the other, despite the fact that predicted values from both modes are not compatible with $Y^{(n)}_{0}$. Therefore, it would seem that Bayes factor model comparisons could be enhanced with our Z-value approach which offers a careful preliminary examination of model adequacy before proceeding to model selection based on Bayes factors.
@@
Figure S20: Thick line: the true BF distributions on samples simulated from $M_{1}$ or $M_{2}$ (depending on the color); Colored area: the estimated BF distributions on samples simulated from $M_{1}$ or $M_{2}$ (depending on the color).